OpenStudy (anonymous):

Estimate the limit numerically lim x→(− Infinity) (x^4 − 4,000x^3)/ (4x^4 + 9,000) ANYONE?! nobody can seem to get this one its not 0 or -1/4 either

myininaya (myininaya):

\[\text{ let } f(x)=\frac{x^4-4000x^3}{4x^4+9000}\] the limit should be 1/4 ( i know this by short cut way)

myininaya (myininaya):

but we need to plug in negative "getting big numbers"

myininaya (myininaya):

\[f(-25)=\frac{(-25)^4-4000(-25)^3}{4(-25)^4+9000}\] make sure when you enter this into the cal you put paranthesis around the negative part you know don't do -25^4 but do (-25)^4 or even (25)^4 since (-25)^4=25^4

myininaya (myininaya):

try it and tell me what you get for x=-25

OpenStudy (anonymous):

yea but its -\[\infty\] haha idk how to do that one on a calculator

myininaya (myininaya):

right

myininaya (myininaya):

you plug in negative "getting large numbers"

myininaya (myininaya):

plug in x=-25 then plug in x=-100 then plug in x=-1000 and tell me what f is getting close to